Determine whether the two lines with the given equations are parallel, perpindicular, or niether with each other
A) -3y=15x-12
B) y=5x+7

Answer by richwmiller(17219)

-3y=15x-12
y=15x/-3-12/-3
y=-5x+4
y=5x+7
-5*5=-25
not parallel and not perpendicular
intersect at x = -3/10, y = 11/2